# Jinhang Chai     Jianqing Fan

### v1

@@ 4 Markov Transition Matrix Estimation
This section showcases the application of our main result to an important problem on the estimation of Markov Transition Matrix. We begin by introducing some necessary notations.
@@ 4.1 Markov Chain Preliminary
Ergodicity is a common assumption for Markov chains. A Markov chain is called irreducible if for each state pair $(i,j)$, there exists a $n_{i,j}\geq 1$ such that $\mathbb{P}(X_{n_{i,j}}=j\big{|}X_{0}=i)>0$. A Markov chain is called aperiodic if for each state $i$, the greatest common divisor of the set $\{n\geq 1\big{|}\mathbb{P}(X_{n}=i\big{|}X_{0}=i)>0\}$ is 1. A Markov chain is ergodic if it is both irreducible and aperiodic. Throughout the paper, we assume $\{X_{0},X_{1},\cdots\}$ is ergodic so there exists a unique stationary(or invariant) distribution $\pi$ which satisfies $\pi P=\pi$ and $\pi\mathbf{1}_{p}=1$. We denote $\pi_{\min}=\min_{i}\pi_{i}$ and $\pi_{\max}=\max_{i}\pi_{i}$.
This paper also uses the frequency matrix $F\in\mathbb{R}^{p\times p}$ which is given by $F_{i,j}=\pi_{i}P_{i,j}$ and represents the probability that a pair of states $(i,j)$ appears in the Markov chain in the long run, that is, $F_{i,j}=\lim_{n\rightarrow\infty}\sum_{k=1}^{n}1_{X_{k}=i,X_{k+1}=j}$. It is straightforward to verify that the frequency matrix has a one-to-one correspondence with the transition matrix. That is, there is a one-to-one correspondence between $\mathcal{F}_{p}$ and $\mathcal{P}_{p}$, where
$\mathcal{F}_{p}:=\{F\in\mathbb{R}^{p\times p}\big{|}\mathbf{1}_{p}^{\top}F\mathbf{1}_{p}=1,F\geq 0\}.$
Reversibility is an important concept. A Markov chain is called reversible if for each pair of states $(i,j)$, $\pi_{i}P_{i,j}=\pi_{j}P_{j,i}$, or equivalently $F_{i,j}=F_{j,i}$. Another important concept in the Markov chain is the mixing time. The $\epsilon$-mixing time is defined as
@@
namely the minimum time that the distribution of the Markov chain is nearly stationary initialized at any value $i$. We also call $\tau^{\star}=\tau(\frac{1}{4})$ the mixing time for short. The mixing time characterizes the speed of the Markov chain converges to the stationary distribution in 1-norm. It has a strong relation with the spectral gap of the transition matrix. Specifically, if the chain is reversible, then we have the following inequality (Levin & Peres,, 2017):
@@ 4.2 Structured Markov Transition Matrix Estimation
Here conditions 1 and 3 in Assumptions 4.1 are standard regularity assumptions in the literature of Markov chain(Norris,, 1998; Chung,, 1967). Ergodicity guarantees that the Markov chain eventually visits each state with enough time which enables good estimation. Technically, it also assures a unique stationary distribution. The second condition is the main structural assumption on the frequency matrix. The third condition requires the stationary distribution not too small to leave some states not well explored (Zhang & Wang,, 2019). We first state a simple result regarding relations between elements of $P$ and $\pi$, whose proof follows from simple algebra and is omitted.
Lemma 4.2.
$P_{\min}\leq\pi_{\min}\leq\pi_{\max}\leq P_{\max},$
@@
$\widetilde{F}_{i,j}=\sum_{k=1}^{n}1_{X_{k}=i,X_{k+1}=j}.$ (8)
In order to apply our deterministic result, we let the observation be $Y=\widetilde{F}$, effective noise be $\widetilde{F}-F$, and apply the incoherent-constrained least-squares (3). To establish its property, we need to bound $\|\widetilde{F}-F^{\star}\|$ as well as $\|\widetilde{F}-F^{\star}\|_{\max}$ with high probability, which is stated in the following proposition. Denote $p_{\max}:=\max_{i,j}p_{i,j}$.
@@
Under Assumption 4.1, for $\forall c_{0}$, there exists a constant $C$ such that for all $n\geq C\tau_{*}p\log^{2}n$, we have
@@
For the max norm error bound, one may attempt to apply the Markov Bernstein Inequality in Jiang et al., (2018) on $Y$. However, it turns out that the absolute spectral gap is 0 for such a particular kind of Markov chain (in Lemma A.1). Hence, this route is infeasible.
@@
The mixing time $\tau_{*}$ in the bound stems from the Markov chain concentration technique (Zhang & Wang,, 2019). Even if the data starts from stationary distribution, in order to decorrelate, traditional concentration inequality is applied individually to groups of samples indexed by arithmetic sequences. As there are $\tau_{*}$ groups of them, there is a loss of factor $\tau_{*}$. The dependence on $\tau_{*}$ makes sense here because even though the chain starts from stationary distribution, there is only a single realized trajectory and the chain has to take $\tau_{*}$ steps to approach stationarity again.
We present our algorithm for Markov transition estimation as follows, which originates from our main method (3).
@@
5:Project $\widehat{F}^{1}$ onto the space of frequency matrices $\mathcal{F}_{p}$ in Euclidean distance(which is equivalent to projection in $L_{1}$ distance) to get $\widehat{F}$. This can done by a rowwise operation. Fix row index $i$, if $e_{i}^{\top}\widehat{F}^{1}\mathbf{1}\leq 1$, we have $\widehat{F}_{i,j}:=\widehat{F}^{1}_{i,j}+\frac{1-e_{i}^{\top}\widehat{F}^{1}\mathbf{1}}{p}$, for every $i,j$. Otherwise, if $e_{i}^{\top}\widehat{F}^{1}\mathbf{1}\geq 1$, find a $t_{i}\geq 0$ such that $\sum_{j=1}^{p}\max\left(\widehat{F}^{1}_{i,j}-t_{i},0\right)=1$ which could be implemented, e.g., by sorting and checking $p$ break points. The projection is then given by $\widehat{F}_{i,j}:=\max\left(\widehat{F}^{1}_{i,j}-t_{i},0\right)$ for every $i,j$.
@@
Now we are ready to present our main theorem in this section. Recall that $p_{\max}:=\max_{i,j}p_{i,j}$ and $\|\cdot\|_{1}$ stands for elementwise $L_{1}$ norm.
@@ 4.3 Implication on structured RL
Specifically, given a (value) vector $v$, we hope to estimate $\mathcal{T}(v):=Pv$, the conditional mean function. The following simple corollary indicates that the low-rank-plus-sparse structure can be utilized to achieve better estimation, in comparison with fully nonparametric estimation.
@@
$\mathbb{E}\|\widehat{\mathcal{T}}(v)-\mathcal{T}(v)\|_{2}^{2}=\operatorname{O}\left(c_{v}\frac{rp}{n}\right).$
@@
From the above corollary, we can see that the structural assumption reduces the estimation error of the expected conditional mean by a factor of ${p}/{r}$. Notably, the proof utilizes the assumption that $v$ is a random variable. The next theorem is devoted to illustrating that even in the exact low-rank case, there is no benefit in the worst case for the deterministic $v$.
In the above corollary, the randomness of $v$ is crucial as the improvement is in the average-case sense. We prove in the following theorem that even in the exact low-rank case where $S^{*}=0$, there is no gain in estimation efficiency in the worst-case(or minimax) sense.
@@
When $v\sim\mathcal{N}(0,I)$, we have $\|u\|_{2}\asymp\|v\|_{2}\asymp p$, therefore, the minimax risk is of order $\operatorname{O}({p^{2}}/{n})$. Compared with Corollary 4.6, it’s larger with a factor of ${p}/{r}$, and there is no gain even assuming the transition matrix is exactly low-rank.
@@
We here provide intuition for the discrepancy between Corollary 4.6 and Theorem 4.7. The estimation error of the conditional mean is a one-dimensional projection of that of the transition matrix estimation. When one of those top singular vectors of $P$ aligns with $v$, the one-dimensional projection estimation error can be large as shown in Theorem 4.7. By randomizing the one-dimensional directions, this unfavorable case of alignment is evened out. This is in analogy to denoising a vector under sparsity condition, where the estimating error in the Euclidean norm is improved by thresholding while the $L_{\infty}$ error typically can’t benefit from the sparsity assumption.
@@ 5.1 Multitask Regression
Apart from the identity observation model, another simple observation model is the multitask regression (Agarwal et al.,, 2012) of the following form:
@@
Let $\Delta_{L}=\widehat{L}-L^{\star},\Delta_{S}=\widehat{S}-S^{\star}$. Assume the design matrix $X$ satisfies that $\sigma_{\max}(X)\leq\sigma_{\max}$, $\sigma_{\min}(X)\geq\sigma_{\min}$, $\|X^{\top}\|_{2,\infty}\leq B$. Under Assumptions 3.1,3.4,3.2, we have
@@ 5.2 Robust Covariance Estimation
In this section, we apply our framework to the problem of covariance estimation, i.e., given data $X_{1},X_{2},\cdots,X_{T}\in\mathbb{R}^{p}$ which follows the same distribution, we aim to estimate the covariance matrix $\Sigma=\mathbb{E}X_{1}X_{1}^{\top}$. A structural assumption on the covariance estimation is that the covariance matrix can be decomposed as the sum of low-rank part and sparse part, i.e.,
@@
$X_{i}=Bf_{i}+\epsilon_{i},\quad f_{i}\in\mathbb{R}^{r}$ (18)
where $f_{i}$ and $\epsilon_{i}$ are uncorrelated.
We consider a general case where we regard $X_{1},X_{2},\cdots,X_{T}$ as an i.i.d. sequence, where $X_{i}$ could potentially have heavy tails. In that case, a natural estimator for general covariance estimation is the Winsorization(truncation) estimator $\widehat{\Sigma}=\{\widehat{\sigma}_{ij}\}_{i,j=1}^{n}$, which is given by
$\widehat{\sigma}_{ij}=\widehat{\mathbb{E}X_{i}X_{j}}^{\top}-\widehat{\mathbb{E}X_{i}}^{\top}\widehat{\mathbb{E}X_{j}}^{\top}$ (19)
@@
$\widehat{\mathbb{E}X_{i}X_{j}}^{\top}$ $=$ $\frac{1}{n}\sum_{i=1}^{\top}\operatorname{sign}(X_{ki}X_{kj})\min(X_{ki}X_{kj},\tau_{1}).$
$\widehat{\mathbb{E}X_{i}}^{\top}$ $=$ $\frac{1}{n}\sum_{i=1}^{\top}\operatorname{sign}(X_{ki})\min(X_{ki},\tau_{2}).$
After the pilot estimator $\widehat{\Sigma}$ is constructed, we apply the optimization program (3) to $\widehat{\Sigma}$ as the observation to get estimators $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$ for low-rank part and sparse part, respectively. The main benefit of using this estimator is that, as shown in the following, the pervasiveness assumption is no longer needed.
We impose the bounded fourth moment assumption common in the literature(Fan et al.,, 2017; Fan et al., 2021a, ; Fan et al., 2021b, ).
@@
Suppose the fourth moment of $X_{i}$ is bounded for every $i$, i.e., $\mathbb{E}(X_{i}^{4})\leq C$.
@@
Denote $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$ as the output of (3). Under Assumptions 3.1,3.2,3.4 on $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$, and Assumption 5.2 on the distribution of $X_{i}$, it holds that
@@
The rate is nearly minimax optimal when $r$ is a constant. In fact, we have $\inf\limits_{\widehat{\Sigma}}\sup\limits_{\Sigma}\|\widehat{\Sigma}-\Sigma\|_{F}^{2}\asymp\frac{p^{2}}{n}$ and $\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}_{F}+\|\widehat{\Sigma}_{S}-\Sigma_{S}\|^{2}_{F}\gtrsim\|\widehat{\Sigma}-\Sigma\|_{F}^{2}$. The inequality $\|W\|\leq p\|W\|_{\max}$ seems loose but is actually the best we can hope for, which is typically the case in factor model when the noise entries are highly dependent, see for instance the example in Fan et al., (2008) where $B=\mathbf{1}_{p}$ and $\operatorname{cov}(\epsilon|f)=I$.
@@
The result shows that we can optimally identify the low-rank part and sparse part separately, with no assumption on the relative magnitude of the low-rank part and sparse part, which is in stark contrast with most existing work on the factor model where pervasiveness is often imposed. It is left to future work to establish the estimation error of the loading vector and in rescaled norm $\|\cdot\|_{\Sigma}$, see details in Fan et al., (2008),Fan et al., (2011),Fan et al., (2013).
When common assumptions in the factor model (18) hold, we can derive the error of estimating loading factors $\{b_{j}\}_{j=1}^{p}$, the proof basically follows Theorem 3 in Fan et al., (2019). Specifically, due to the identifiability issue, we can assume $cov(f_{i})=I$ and columns of $B$ are pairwise orthogonal (Fan et al.,, 2013). Denote $l^{th}$ column of $B$ as $\widetilde{b}_{l}$. We know that $v_{l}=\frac{\widetilde{b}_{l}}{\|\widetilde{b}_{l}\|}$ and $\lambda_{l}=\|\widetilde{b}_{l}\|^{2}$ are singular vectors and singular values of $B$, respectively. Further, $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{r}$ are in descending order. Now let $\widehat{v}_{l}$ and $\widehat{\lambda}_{l}$ be the singular vectors and singular values of $\widehat{\Sigma}_{L}$, respectively. We get an estimator of $B$ as $\widehat{B}=[\widehat{\lambda}_{1}^{\frac{1}{2}}\widehat{v}_{1},\cdots,\widehat{\lambda}_{r}^{\frac{1}{2}}\widehat{v}_{r}]$. Denote the $j^{th}$ row of $\widehat{B}$ as $\widehat{b}_{j}$. The following theorem gives the estimation error of $\widehat{b}_{j}$.
Theorem 5.4.
Under the same assumptions of Theorem 5.3, and further assuming that $\lambda_{l}-\lambda_{l+1}=\Omega(p)$ for all $1\leq l\leq r$ and $n\gtrsim\frac{p\log^{2}p}{\mu}$, we have
$\min_{H\in{\cal O}^{r\times r}:H^{2}=\mathbf{I}}\sum_{j=1}^{p}\|b_{j}-H\widehat{b}_{j}\|^{2}\lesssim\frac{r^{3}p\log^{2}p}{n}.$

### v2

@@ 3 Markov Transition Matrix Estimation
This section showcases the application of our main result to an important problem in the estimation of the Markov Transition Matrix. We begin by introducing some necessary notations.
@@ 3.1 Markov Chain Preliminary
Ergodicity is a common assumption for Markov chains. A Markov chain is called irreducible if, for each state pair $(i,j)$, there exists a $n_{i,j}\geq 1$ such that $\mathbb{P}(X_{n_{i,j}}=j\big{|}X_{0}=i)>0$. A Markov chain is called aperiodic if for each state $i$, the greatest common divisor of the set $\{n\geq 1\big{|}\mathbb{P}(X_{n}=i\big{|}X_{0}=i)>0\}$ is 1. A Markov chain is ergodic if it is both irreducible and aperiodic. Throughout the paper, we assume $\{X_{0},X_{1},\cdots\}$ is ergodic so there exists a unique stationary(or invariant) distribution $\pi$ which satisfies $\pi P=\pi$ and $\pi\mathbf{1}_{p}=1$. We denote $\pi_{\min}=\min_{i}\pi_{i}$ and $\pi_{\max}=\max_{i}\pi_{i}$. Ergodicity implies $\pi_{\min}>0$.
This paper also uses the frequency matrix $F\in\mathbb{R}^{p\times p}$ which is given by $F_{i,j}=\pi_{i}P_{i,j}$ and represents the probability that a pair of states $(i,j)$ appears in the Markov chain in the long run, that is, $F_{i,j}=\lim_{n\rightarrow\infty}(1/n)\sum_{k=1}^{n}1_{X_{k}=i,X_{k+1}=j}$. To facilitate theoretical analysis, we define the enlarged frequency matrix class as
$\mathcal{F}_{p}:=\{F\in\mathbb{R}^{p\times p}\big{|}\mathbf{1}_{p}^{\top}F\mathbf{1}_{p}=1,\ F\geq 0,\ e_{i}^{\top}F\mathbf{1}_{p}>0,\ \forall i\}.$ (9)
Note that a proper frequency matrix further needs to satisfy $\sum_{i=1}^{p}F_{i,k}=\pi_{k}=\sum_{j=1}^{p}F_{k,j}$. By incorporating these constraints, the proper frequency matrix class ~$\tilde{{\cal F}}_{p}$, as a nontrivial subset of ${\cal F}_{p}$, can be defined as
~$\tilde{\mathcal{F}}_{p}:=\{F\in{\cal F}_{p}\big{|}e_{i}^{\top}F\mathbf{1}_{p}=\mathbf{1}_{p}^{\top}Fe_{i},\ \forall i\}.$ (10)
One can verify that every $F\in\tilde{{\cal F}}_{p}$ is a proper frequency matrix. In fact, by defining $\pi=e_{i}^{\top}F\mathbf{1}_{p}$ and $P$ as $P_{i,j}=\frac{e_{i}^{\top}Fe_{j}}{e_{i}^{\top}F\mathbf{1}_{p}}$, we can argue that $\pi$ is the stationary distribution and $P$ is a valid transition matrix, leading to $F_{i,j}=P_{i,j}\pi_{i}$.
Similarly, we can argue that any $F\in{\cal F}_{p}$ corresponds to an induced transition matrix $P$. On the other hand, any $P$ in ${\cal P}_{p}$ representing ergodic chain corresponds to an induced frequency matrix $F\in\tilde{{\cal F}}_{p}$.
Reversibility is another important concept. A Markov chain is called reversible if for each pair of states $(i,j)$, $\pi_{i}P_{i,j}=\pi_{j}P_{j,i}$, or equivalently $F_{i,j}=F_{j,i}$. Another important concept in the Markov chain is the mixing time. The $\epsilon$-mixing time is defined as
@@
namely the minimum time that the distribution of the Markov chain is nearly stationary initialized at any value $i$. We also call $\tau_{*}=\tau(\frac{1}{4})$ the mixing time for short. The mixing time characterizes the speed of the Markov chain converges to the stationary distribution in 1-norm. It has a strong relation with the spectral gap of the transition matrix. Specifically, if the chain is reversible, then we have the following inequality (Levin & Peres,, 2017):
@@ 3.2 Structured Markov Transition Matrix Estimation
Here, conditions 1 and 3 in Assumptions 3.1 are standard regularity assumptions in the literature of Markov chain (Norris,, 1998; Chung,, 1967). Ergodicity guarantees that the Markov chain eventually visits each state with enough time which enables good estimation. Technically, it also assures a unique stationary distribution. The second condition is the main structural assumption on the frequency matrix. The third condition ensures that the probability of each state in the stationary distribution is sufficiently large, preventing some states from being inadequately explored (Zhang & Wang,, 2019).
We provide additional illustrations to clarify the second condition. While Zhang & Wang, (2019) assumes a low-rank structure for the transition matrix, such as in the case of state aggregation, this assumption is often overly restrictive. In practice, even if states can be grouped, the interactions among groups are not always homogeneous. Moreover, certain pairs of nodes may exhibit stronger or weaker connections, leading to deviations from exact low-rankness. In principle, the low-rank part may capture systematic trends, while the sparse part characterizes anomalies or unusual events. Below, we present several concrete examples to demonstrate the idea.
  1. 1.
    Population Flow: In urban planning, ride-sharing services, or epidemiology, one may care about how people move between different regions. The low-rank part can represent the general movement trends, such as people commuting daily between residential and work areas. On the other hand, the sparse part can capture rare events, such as movement spikes due to festivals and emergencies or unusual travel patterns during holidays.
  2. 2.
    Financial Markets: One can model the transition of financial instruments between investors using Markov chains. In this example, the low-rank part may capture typical trading patterns, such as systematic trading strategies that the investor deploys. On the other hand, the sparse part may represent rare events like market crashes or unexpected news affecting a single stock or sector.
  3. 3.
    Social Network Interactions: Consider modeling interactions (e.g., messages, likes, or posts) between users in a social network. In this example, the low-rank component reflects regular interactions, such as users engaging primarily with close friends or family. On the other hand, sparse component captures irregular interactions, such as viral content, causing engagement spikes with unfamiliar users.
Example.
As an illustrative example, consider a Markov chain modeling population flow across four locations, ${\cal S}=\{A,B,C,D\}$. Under normal conditions, one may impose the state aggregation structure (Definition 3 in Zhang & Wang, (2019)) with $\Omega_{1}=\{A,B\}$ and $\Omega_{2}=\{C,D\}$, resulting in the low-rank structure. Suppose the transition matrix under normal conditions is given by
[…]
However, there may be aspects not fully captured by the state aggregation model. For example, suppose during some time period, the transportation from $D$ to other three places is under construction, while the transportation from other places to $D$ remains unchanged. In this case, the true transition matrix in this period is given by
[…]
Therefore, $P^{\star}$ has the low-rank-plus-sparse structure, and the same holds for $F^{\star}=\operatorname{diag}(\pi^{\star})P^{\star}$.∎
@@
$\widetilde{F}_{i,j}=(1/n)\sum_{k=1}^{n}1_{X_{k}=i,X_{k+1}=j}.$ (11)
In order to apply our deterministic result, we let the observation be $Y=\widetilde{F}$, effective noise be $\widetilde{F}-F$, and apply the incoherent-constrained least-squares (3). To establish its property, we need to bound $\|\widetilde{F}-F^{\star}\|$ as well as $\|\widetilde{F}-F^{\star}\|_{\max}$ with high probability, which is stated in the following proposition. In particular, the first half of the proposition is the same as in Lemma 7 in Zhang & Wang, (2019). Denote $p_{\max}:=\max_{i,j}p_{i,j}$.
@@
Under Assumption 3.1, for $\forall c_{0}$, there exists a constant $C$ such that for all $n\geq C\tau_{*}p\log^{2}n$, we have [Lemma 7 of Zhang & Wang, (2019)]
@@
For the max norm error bound, one may attempt to apply Bernstein’s inequality for Markov chain in Jiang et al., (2018) on $Y$. However, it turns out that the absolute spectral gap is zero for this specific type of Markov chain (as shown in Lemma B.1 in the supplementary material), making this approach infeasible.
@@
The mixing time $\tau_{*}$ in the bound stems from the Markov chain concentration technique (Zhang & Wang,, 2019). Even if the data starts from stationary distribution, in order to decorrelate, traditional concentration inequality is applied individually to groups of samples indexed by arithmetic sequences. As there are $\tau_{*}$ groups of them, there is a loss of factor $\tau_{*}$. The dependence on $\tau_{*}$ makes sense here because even though the chain starts from stationary distribution, there is only a single realized trajectory, and the chain has to take $\tau_{*}$ steps to approach stationarity again.
We present our algorithm for Markov transition estimation in detail in Algorithm 2, which originates from our main method (3). After solving (3), we project the solution to ${\cal F}_{p}$ and then normalize each row to obtain the estimated transition matrix $\widehat{P}$. It is worth noting that while $\widehat{P}$ is a valid transition matrix, $\widehat{F}$ may not be a valid frequency matrix; see (9) and discussions therein. One could either project the solution to the proper frequency class ~$\tilde{{\cal F}}_{p}$ defined in (10), which forms a polytope, or transform $\widehat{P}$ back. The technical details are omitted for brevity.
@@
5:Project $\widehat{F}^{1}$ onto the enlarged space of frequency matrices $\mathcal{F}_{p}$ in Euclidean distance(which is equivalent to projection in $L_{1}$ distance) to get $\widehat{F}$. This can be done by the following. Find a $t$ such that $\sum_{i=1}^{p}\sum_{j=1}^{p}\max\left(\widehat{F}^{1}_{i,j}-t,0\right)=1$, which could be implemented, e.g., by sorting and checking $p$ breakpoints or use bisection method. The projection is then given by $\widehat{F}_{i,j}:=\max\left(\widehat{F}^{1}_{i,j}-t,0\right)$.
@@
Now, we are ready to present our main theorem in this section. Recall that $p_{\max}:=\max_{i,j}p_{i,j}$ and $\|\cdot\|_{1}$ stands for elementwise $L_{1}$ norm.
@@
Remark 8.
The rate with respect to sample size $n$ scales as $\sqrt{\log^{2}n/n}$. The logarithmic terms arise from applying the union bound. While we generally disregard the logarithmic terms in this paper, it is an interesting question whether they can be further eliminated. Another interesting question is whether multi-cover time instead of mixing time can be used to achieve tighter bounds (Chan et al.,, 2021).
@@ 3.3 Implication on structured RL
Specifically, given a (value) vector $v$, we hope to estimate $\mathcal{T}(v):=Pv$, the conditional mean function. The following simple corollary indicates that the low-rank-plus-sparse structure can be utilized to achieve better estimation in comparison with fully nonparametric estimation.
@@
$\mathbb{E}\|\widehat{\mathcal{T}}(v)-\mathcal{T}(v)\|_{2}^{2}=\operatorname{O}\left(c_{v}\frac{rp\log^{2}n}{n}\right).$
@@
From the above corollary, we observe that the structural assumption (low-rank-plus-sparse) reduces the estimation error of the expected conditional mean by a factor of ${p}/{r}$compared to the scenario without such an assumption. Notably, the proof utilizes the assumption that $v$ is a random variable. The next theorem is devoted to illustrating that even in the exact low-rank case, there is no benefit in the worst case for the deterministic $v$.
In the above corollary, the randomness of $v$ is crucial as the improvement is in the average-case sense. We prove in the following theorem that even in the exact low-rank case where $S^{\star}=0$, there is no gain in estimation efficiency in the worst-case(or minimax) sense.
@@
When $v\sim\mathcal{N}(0,I)$, we have $\|u\|_{2}\asymp\|v\|_{2}\asymp p$. Therefore, the minimax risk is of order $\operatorname{O}({p^{2}}/{n})$. Compared with Corollary 3.5, it’s larger with a factor of ${p}/{r}$, and there is no gain even assuming the transition matrix is exactly low-rank.
@@
We here provide intuition for the discrepancy between Corollary 3.5 and Theorem 3.6. The estimation error of the conditional mean is a one-dimensional projection of that of the transition matrix estimation. When one of those top singular vectors of $P$ aligns with $v$, the one-dimensional projection estimation error can be large, as shown in Theorem 3.6. By randomizing the one-dimensional directions, this unfavorable case of alignment is evened out. This is in analogy to denoising a vector under sparsity condition, where the estimating error in the Euclidean norm is improved by thresholding while the $L_{\infty}$ error typically can’t benefit from the sparsity assumption.
@@ 4.1 Multitask Regression
Apart from the identity observation model, another simple observation model is the structured multitask regression (Ando et al.,, 2005; Argyriou et al.,, 2008; Kumar & Daume III,, 2012). In particular, Agarwal et al., (2012) proposed the following low-rank-plus-sparse form:
@@
Let $\Delta_{L}=\widehat{L}-L^{\star},\Delta_{S}=\widehat{S}-S^{\star}$. Assume the design matrix $X$ satisfies that $\sigma_{\max}(X)\leq\sigma_{\max}$, $\sigma_{\min}(X)\geq\sigma_{\min}$, $\|X^{\top}\|_{2,\infty}\leq B$. Under Assumptions 2.1,2.2,2.4, we have
@@
Remark 14.
The Corollary 4 in Agarwal et al., (2012) gives a similar deterministic result, albeit with a trailing term ${\cal O}(s/d^{2})$ that doesn’t disappear even $n\rightarrow\infty$. When $s=0$, the problem simplifies to low-rank multi-task learning or reduced-rank regression, and similar probabilistic results can be found in Duan & Wang, (2023). Multi-task learning is also closely related to trace regression (Anderson,, 1951; Izenman,, 1975).
@@ 4.2 Robust Covariance Estimation
In this section, we apply our framework to the problem of covariance estimation (Fan et al., 2019b, ), i.e., given data $X_{1},X_{2},\cdots,X_{T}\in\mathbb{R}^{p}$ which follows the same distribution, we aim to estimate the covariance matrix $\Sigma=\mathbb{E}X_{1}X_{1}^{\top}$. A structural assumption on the covariance estimation is that the covariance matrix can be decomposed as the sum of low-rank part and sparse part, i.e.,
@@
$X_{t}=Bf_{t}+\epsilon_{t},\quad f_{t}\in\mathbb{R}^{r}$ (20)
where $f_{t}$ and $\epsilon_{t}$ are uncorrelated and $\operatorname{var}(\epsilon_{t})=\Sigma_{S}$.
We consider a general case where we regard $X_{1},X_{2},\cdots,X_{T}$ as an i.i.d. sequence, where $X_{t}$ could potentially have heavy tails. In that case, a natural estimator for general covariance estimation is the Winsorization (truncation) estimator $\widehat{\Sigma}=\{\widehat{\sigma}_{ij}\}_{i,j=1}^{p}$, whose entries are given by [^7] For notational simplicity, let $X_{ki}$ denote the $i$-th entry of $X_{k}$.
$\widehat{\sigma}_{ij}=\widehat{\mathbb{E}X_{\cdot i}X_{\cdot j}}-\widehat{\mathbb{E}X_{\cdot i}}\widehat{\mathbb{E}X_{\cdot j}},\quad i,j=1,2,\cdots,p$ (21)
@@
$\widehat{\mathbb{E}X_{\cdot i}X_{\cdot j}}$ $=$ $\frac{1}{n}\sum_{k=1}^{T}\operatorname{sign}(X_{ki}X_{kj})\min(|X_{ki}X_{kj}|,\tau_{1}).$
$\widehat{\mathbb{E}X_{\cdot i}}$ $=$ $\frac{1}{n}\sum_{k=1}^{T}\operatorname{sign}(X_{ki})\min(|X_{ki}|,\tau_{2}).$
After the pilot estimator $\widehat{\Sigma}$ is constructed, we apply the optimization program (3) to $\widehat{\Sigma}$ as the observation to get estimators $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$ for low-rank part and sparse part, respectively. The main benefit of using this estimator is that, as shown in the following, the pervasiveness assumption is no longer needed. That is an advantage of the optimization-based method.
We impose the bounded fourth-moment assumption common in the literature (Fan et al.,, 2017; Fan et al., 2021a, ; Fan et al., 2021b, ).
@@
Suppose the fourth moment of $X_{1i}$ is bounded for every $i$, i.e., $\mathbb{E}(X_{1i}^{4})\leq C$.
@@
Denote $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$ as the output of (3). Under Assumptions 2.1,2.2,2.4 on $\widehat{\Sigma}_{L}$ and $\widehat{\Sigma}_{S}$, and Assumption 4.2 on the distribution of $X_{i}$, it holds that
@@
The rate is nearly minimax optimal when $r$ is a constant. In fact, we have $\inf\limits_{\widehat{\Sigma}}\sup\limits_{\Sigma}\|\widehat{\Sigma}-\Sigma\|_{F}^{2}\asymp\frac{p^{2}}{n}$ and $\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}_{F}+\|\widehat{\Sigma}_{S}-\Sigma_{S}\|^{2}_{F}\gtrsim\|\widehat{\Sigma}-\Sigma\|_{F}^{2}$.
When common assumptions on the factor model (20) hold, we can derive the error of estimating loading factors $\{b_{j}\}_{j=1}^{p}$, and the proof basically follows Theorem 3 in Fan et al., 2019a . Specifically, due to the identifiability issue, we can assume $cov(f_{i})=I$ and columns of $B$ are pairwise orthogonal (Fan et al.,, 2013). Denote $l^{th}$ column of $B$ as $\widetilde{b}_{l}$. We know that $v_{l}=\frac{\widetilde{b}_{l}}{\|\widetilde{b}_{l}\|}$ and $\lambda_{l}=\|\widetilde{b}_{l}\|^{2}>0$ are singular vectors and singular values of $B$, respectively. Further, $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{r}$ are in descending order. Now let $\widehat{v}_{l}$ and $\widehat{\lambda}_{l}$ be the singular vectors and singular values of $\widehat{\Sigma}_{L}$, respectively. We get an estimator of $B$ as $\widehat{B}=[\widehat{\lambda}_{1}^{\frac{1}{2}}\widehat{v}_{1},\cdots,\widehat{\lambda}_{r}^{\frac{1}{2}}\widehat{v}_{r}]$. Denote the $j^{th}$ row of $\widehat{B}$ as $\widehat{b}_{j}$. The following theorem gives the estimation error of $\widehat{b}_{j}$.
Theorem 4.4.
Under the same assumptions of Theorem 4.3, we have with probability at least $1-n^{-c}$ for some constant $c>0$,
$\min_{H\in\mathbb{R}^{r\times r}}\|BH-\widehat{B}\|^{2}\lesssim\frac{rp^{2}\log^{2}(p)\log n}{n\lambda_{r}}.$ (22)
Furthermore, if $\lambda_{l}-\lambda_{l+1}=\Omega(p)$ for all $1\leq l\leq r$ and $n\gtrsim\frac{p\log^{2}p}{\mu}$, we have with probability at least $1-n^{-c}$,
$\min_{H\in{\cal O}^{r\times r}:H^{2}=\mathbf{I}}\sum_{j=1}^{p}\|b_{j}-H\widehat{b}_{j}\|^{2}\lesssim\frac{r^{3}p\log^{2}(p)\log n}{n}.$ (23)
@@
Note that there is rotational ambiguity regarding the direction factors. Hence, an orthogonal (or nearly orthogonal) transformation is used to eliminate this ambiguity, which in (23) is taken to be a diagonal matrix with ±$\pm 1$ on the diagonal.